# Cohomology and Deformations of left-symmetric Rinehart Algebras

### v3

@@ 5 The Matsushima-Nijenhuis bracket for left-symmetric Rinehart algebras
In this section we construct a graded Lie algebra whose Maurer-Cartan elements are left-symmetric Rinehart algebras which gives a rise to a coboundary operator.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebras. A multiderivation of degree $n$ is a multilinear map $P\in\mathrm{Hom}(\Lambda^{n}L\otimes L,L))$ such that for any $a\in A$ and $x_{i}\in L,\ i=1,2,\cdots n+1$, we have
$P(x_{1},\dots,ax_{i},\dots,x_{n},x_{n+1})=aP(x_{1},\dots,x_{i},\dots,x_{n},x_{n+1}),\ i=1,\cdots,n,$ (5.1)
$P(x_{1},\cdots,x_{n},ax_{n+1})=aP(x_{1},\cdots,x_{n},x_{n+1})+\Xi_{P}(x_{1},\cdots,x_{n})(a)x_{n+1},$ (5.2)
@@
A permutation $\sigma\in\mathbb{S}_{n}$ is called an $(i,n-i)$-unshuffle if $\sigma(1)<\cdots<\sigma(i)$ and $\sigma(i+1)<\cdots<\sigma(n)$. If $i=0$ and $i=n$, we assume $\sigma=Id$. The set of all $(i,n-i)$-unshuffles will be denoted by $\mathbb{S}_{(i,n-i)}$. The notion of an $(i_{1},\cdots,i_{k})$-unshuffle and the set $\mathbb{S}_{(i_{1},\cdots,i_{k})}$ are defined similarly.
Let $P\in\mathfrak{D}^{m}(L)$ and $Q\in\mathfrak{D}^{n}(L)$, we define the Matsushima-Nijenhuis bracket $[\cdot,\cdot]_{MN}:\mathfrak{D}^{m}(L)\times\mathfrak{D}^{n}(L)\to\mathfrak{D}^{m+n}(L)$ by
$[P,Q]_{MN}=P\diamond Q-(-1)^{mn}Q\diamond P,$ (5.3)
@@
$+(-1)^{mn}\sum_{\sigma\in\mathbb{S}_{(n,m)}}(-1)^{\sigma}P(x_{\sigma(1)},\cdots,x_{\sigma(n)},Q(x_{\sigma(n+1)},x_{\sigma(n+2)},\cdots,x_{\sigma(m+n)},x_{m+n+1})).$
@@
It is straightaorward to check that $[P,Q]_{MN}$ is skew-symmetric with respect to its first $m+n$ arguments. Thus $[P,Q]_{MN}$ is $A$-linear with respect to its first $m+n$ arguments.
On the other hand, following to a straightforward calculation, we have
@@
$\delta(P)=(-1)^{n-1}[\pi,P]_{MN},\quad\forall P\in\mathfrak{D}^{n-1}(L).$ (5.4)
@@
$\Xi_{P}(x_{1},x_{2},\cdots,x_{n})$ $=$ $\sum_{i=1}^{n}(-1)^{i+1}[\Xi_{\pi}(x_{i}),\Xi_{P}(x_{1},x_{2},\cdots,\hat{x_{i}},\cdots,x_{n})]$ (5.6)
@@ 6 Deformation of left-symmetric Rinehart algebra
We investigate in this section a deformation theory of a left-symmetric Rinehart algebra. But first let introduce some notations. For a left-symmetric Rinehart algebra $(L,A,\mathfrak{m},\ell)$ we will denote the left-symmetric multiplication """"$\cdot$"" ⋅ " by $\mathfrak{m}$ in the sequel of the paper. Let $\mathbb{K}[[t]]$ be the formal power series ring in one variable $t$ and coefficients in $\mathbb{K}$. Let $L[[t]]$ be the set of formal power series whose coefficients are elements of $L$ (note that $L[[t]]$ is obtained by extending the coefficients domain of $L$ from $\mathbb{K}$ to $\mathbb{K}[[t]]$). Then $L[[t]]$ is $\mathbb{K}[[t]]$-module.
@@ 6.1 Formal deformations
[…]
@@
Moreover, since $[\mathfrak{m}_{t},\mathfrak{m}_{t}]_{MN}=0$, then it corresponds to a left symmetric Lie Rinehart algebra structure. In particular, it yields a $t$-parametrized family of product
$\mathfrak{m}_{t}:L\otimes L\to L$
and a familly of maps
$\ell_{t}:L\rightarrow Der(A)$
which satisfy the following identities:
@@
$\ell_{t}(x)=$ $\ell(x)+\displaystyle\sum_{i\geq 1}t^{i}\Xi_{\mathfrak{m}_{i}}(x)$
for all $x,y\in L$. The $t$-parametrized family $(L,A,\mathfrak{m}_{t},\ell_{t})$ is called a $1$-parameter formal deformation of $(L,A,\mathfrak{m},\ell)$ generated by $\mathfrak{m}_{1},\cdots,\mathfrak{m}_{m}\in\mathfrak{D}^{1}(L)$.
@@
$\mathfrak{m}_{t}(\mathfrak{m}_{t}(x,y),z)-\mathfrak{m}_{t}(x,\mathfrak{m}_{t}(y,z))=\mathfrak{m}_{t}(\mathfrak{m}_{t}(y,x),z)-\mathfrak{m}_{t}(y,\mathfrak{m}_{t}(x,z)).$ (6.1)
$\mathfrak{m}_{t}(ax,y)=a\mathfrak{m}_{t}(x,y),$ (6.2)
$\mathfrak{m}_{t}(x,ay)=a\mathfrak{m}_{t}(x,y)+\ell_{t}(x)ay$ (6.3)
The identities (6.2)-(6.3), mean that $\mathfrak{m}_{i}\in\mathfrak{D}^{1}(L)$. Comparing the coefficients of $t^{n}$ for $n\geq 0$ in Eq.(6.1), we get the following:
$\sum_{i+j=n}\mathfrak{m}_{i}(\mathfrak{m}_{j}(x,y),z)-\mathfrak{m}_{i}(x,\mathfrak{m}_{j}(y,z))-\mathfrak{m}_{i}(\mathfrak{m}_{j}(y,x),z)+\mathfrak{m}_{i}(y,\mathfrak{m}_{j}(x,z))=0.$ (6.4)
For $n=1$, Eq. (6.4) implies
@@
The $1$-degree multiderivation $\mathfrak{m}_{1}$ is called the infinitesimal of the deformation $\mathfrak{m}_{t}$. More generally, if $m_{i}=0$ for $11\leq i\leq n-$ 1 and $\mathfrak{m}_{n}$ is non zero $1$-degree multiderivation then $\mathfrak{m}_{n}$ is called the $n$-infinitesimal of the deformation $\mathfrak{m}_{t}$. By the above discussion, the following proposition follows immediately.
@@
Now we give a notion of equivalence of two deformations. Let us denote a deformation $(L,A,\mathfrak{m}_{t},\ell_{t})$ of $(L,A,\mathfrak{m},\ell)$ simply by $L_{t}$. Let us consider two deformations $L_{t}$ and $L^{\prime}_{t}=(L,A,\mathfrak{m}^{\prime}_{t},\ell^{\prime}_{t})$ of $(L,A,\mathfrak{m},\ell)$, generated by $\mathfrak{m}_{i}$ and $\mathfrak{m}^{\prime}_{i}$, respectively for $i\geq 0$.
@@
$\Phi_{t}:L[[t]]\rightarrow L[[t]]~{}~{}\mbox{ defined as}~{}~{}\Phi_{t}=id_{L}+\sum_{i\geq 1}t^{i}\phi_{i}$
@@
Proof. Let $\Phi_{t}$ be an equivalence of deformation between $\mathfrak{m}_{t}$ and $\tilde{\mathfrak{m}}_{t}$. Then we get,
@@
$\tilde{\mathfrak{m}}_{1}(x,y)+\Phi_{1}(\mathfrak{m}_{0}(x,y))=\mathfrak{m}_{1}(x,y)+\mathfrak{m}_{0}(\Phi_{1}(x),y)+\mathfrak{m}_{0}(x,\Phi_{1}(y))$
@@
So, cohomology class of infinitesimal of the deformation is determined by the equivalence class of deformation of $\mathfrak{m}_{t}$.
@@
Proof. Let $\mathfrak{m}_{t}$ be a deformation of left-symmetric Rinehart algebra with n-infinitesimal $m_{n}$ for some $n\geq 1$. Assume that there exists a 2-cochain $\phi\in C^{1}(L,L)$ with $\delta(\phi)=m_{n}$. Then set
@@
So, $\bar{m_{n}}=0$. We can repeat the argument to kill off any infinitesimal, which is a coboundary.
@@ 6.2 Obstructions to the extension theory of deformations
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra. Now we consider the problem of extending a deformation of $\mathfrak{m}$ of order $n$ to a deformation of $\mathfrak{m}$ of order $(n+1)$. Let $m_{t}$ and $\ell_{t}$ be a deformation of order $n$ of $\mathfrak{m}$ and $\ell$ respectively. That is
@@
where $m_{i}\in\mathfrak{D}^{1}(L)$ and $\ell_{i}:L\to Der(A)$ a linear map for each $1\leq i\leq n$ such that
$\mathfrak{m}_{i}(\mathfrak{m}_{j}(x,y),z)-\mathfrak{m}_{i}(x,\mathfrak{m}_{j}(y,z))=\mathfrak{m}_{i}(\mathfrak{m}_{j}(y,x),z)-\mathfrak{m}_{i}(y,\mathfrak{m}_{j}(x,z)),$ (6.5)
$\mathfrak{m}_{i}(ax,y)=a\mathfrak{m}_{i}(x,y),$ (6.6)
$\mathfrak{m}_{i}(x,ay)=a\mathfrak{m}_{i}(x,y)+\ell_{i}(x)ay$ (6.7)
for all $1\leq i,j\leq n$. If there exists a 2-cochain $m_{n+1}\in\mathfrak{D}^{1}(L)$ and $\ell_{n+1}:L\to Der(A)$ such that $(L,A,\tilde{\mathfrak{m}_{t}},\tilde{\ell_{t}})$ is a deformation of $(L,A,\mathfrak{m},\ell)$ of order $n+1$, where
@@
Let $\mathfrak{m}_{t}$ be a deformation of $\mathfrak{m}$ of order $n$. Consider the cochain $Obs_{L}\in C^{3}(L,L)$, where
$Obs_{L}(x,y,z)=\sum\limits_{i+j=n+1;~{}i,j>0}\Big{(}\mathfrak{m}_{i}(\mathfrak{m}_{j}(x,y),z)-\mathfrak{m}_{i}(x,\mathfrak{m}_{j}(y,z))-\mathfrak{m}_{i}(\mathfrak{m}_{j}(y,x),z)+\mathfrak{m}_{i}(y,\mathfrak{m}_{j}(x,z)))\Big{)},$ (6.8)
@@
Proof. Suppose that a deformation $\mathfrak{m}_{t}$ of order $n$ extends to a deformation of order $n+1$. Then
$\sum_{i+j=n+1;~{}i,j\geq 0}\Big{(}\mathfrak{m}_{i}(\mathfrak{m}_{j}(x,y),z)-\mathfrak{m}_{i}(x,\mathfrak{m}_{j}(y,z))-\mathfrak{m}_{i}(\mathfrak{m}_{j}(y,x),z)+\mathfrak{m}_{i}(y,\mathfrak{m}_{j}(x,z)))\Big{)}=0.$
@@
$Obs_{L}=\delta(m_{n+1})$
for some 2-cochain $m_{n+1}$. Define a map $\tilde{\mathfrak{m}_{t}}:L[[t]]\times L[[t]]\rightarrow L[[t]]$ as follows
$\tilde{\mathfrak{m}_{t}}=\mathfrak{m}_{t}+m_{n+1}t^{n+1}.$
@@
$\sum_{i+j=n+1;~{}i,j\geq 0}\Big{(}\mathfrak{m}_{i}(\mathfrak{m}_{j}(x,y),z)-\mathfrak{m}_{i}(x,\mathfrak{m}_{j}(y,z))-\mathfrak{m}_{i}(\mathfrak{m}_{j}(y,x),z)+\mathfrak{m}_{i}(y,\mathfrak{m}_{j}(x,z)))\Big{)}=0.$
@@
If $H^{3}(L,L)=0,$ then every 2-cocycle in $C^{2}(L,L)$ is an infinitesimal of some deformation of $\mathfrak{m}$.

### v5

@@ 5 The Matsushima-Nijenhuis bracket for left-symmetric Rinehart algebras
In this section we construct a graded Lie algebra whose Maurer-Cartan elements are left-symmetric Rinehart algebras which give rise to a coboundary operator.
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebras. A multiderivation of degree $n$ is a multilinear map $P\in\mathrm{Hom}(\Lambda^{n}L\otimes L,L)$ such that, for every $a\in A$, $x_{i}\in L$ and $i\in\{1,2,\dotsc,n+1\}$, we have
$P(x_{1},\dots,ax_{i},\dots,x_{n},x_{n+1})=aP(x_{1},\dots,x_{i},\dots,x_{n},x_{n+1}),$ (11)
$P(x_{1},\cdots,x_{n},ax_{n+1})=aP(x_{1},\cdots,x_{n},x_{n+1})+\Xi_{P}(x_{1},\cdots,x_{n})(a)x_{n+1},$ (12)
@@
A permutation $\sigma\in\mathbb{S}_{n}$ is called an $(i,n-i)$-unshuffle if $\sigma(1)<\cdots<\sigma(i)$ and $\sigma(i+~{}1)<~{}\cdots<\sigma(n)$. If $i=0$ and $i=n$, we assume $\sigma=Id$. The set of all $(i,n-i)$-unshuffles will be denoted by $\mathbb{S}_{(i,n-i)}$. The notion of an $(i_{1},\cdots,i_{k})$-unshuffle and the set $\mathbb{S}_{(i_{1},\cdots,i_{k})}$ are defined similarly.
Let $P\in\mathfrak{D}^{m}(L)$ and $Q\in\mathfrak{D}^{n}(L)$. We define the Matsushima–Nijenhuis bracket $[\cdot,\cdot]_{MN}:~{}\mathfrak{D}^{m}(L)\times\mathfrak{D}^{n}(L)\to\mathfrak{D}^{m+n}(L)$ by
$[P,Q]_{MN}=P\diamond Q-(-1)^{mn}Q\diamond P,$
@@
$\quad+(-1)^{mn}\sum_{\sigma\in\mathbb{S}_{(n,m)}}(-1)^{\sigma}P(x_{\sigma(1)},\cdots,x_{\sigma(n)},Q(x_{\sigma(n+1)},x_{\sigma(n+2)},\cdots,x_{\sigma(m+n)},x_{m+n+1})).$
@@
It is straightforward to check that $[P,Q]_{MN}$ is skew-symmetric with respect to its first $m+n$ arguments. Thus $[P,Q]_{MN}$ is $A$-linear with respect to its first $m+n$ arguments.
On the other hand, following a straightforward calculation, we have
@@
$\delta(P)=(-1)^{n-1}[\pi,P]_{MN},\quad\forall P\in\mathfrak{D}^{n-1}(L).$ (13)
@@
$\Xi_{P}(x_{1},x_{2},\cdots,x_{n})$ $=$ $\sum_{i=1}^{n}(-1)^{i+1}[\Xi_{\pi}(x_{i}),\Xi_{P}(x_{1},x_{2},\cdots,\hat{x_{i}},\cdots,x_{n})]$ (15)
@@ 6 Deformation of left-symmetric Rinehart algebra
We investigate in this section a deformation theory of left-symmetric Rinehart algebras. But first let us introduce some notation. For a left-symmetric Rinehart algebra $(L,A,\mathfrak{m},\ell)$ we will denote the left-symmetric multiplication ``"``"`$\cdot$"` ` ⋅ " by $\mathfrak{m}$ in the sequel of the paper. Let $\mathbb{K}[[t]]$ be the formal power series ring in one variable $t$ and coefficients in $\mathbb{K}$. Let $L[[t]]$ be the set of formal power series whose coefficients are elements of $L$ (note that $L[[t]]$ is obtained by extending the coefficients domain of $\mathbb{K}[[t]]$ from $\mathbb{K}$ to $L$). Thus, $L[[t]]$ is a $\mathbb{K}[[t]]$-module.
@@ 6.1 Formal deformations
which is given by $\mathfrak{m}_{t}(x,y)=\displaystyle\sum_{i\geq 0}t^{i}\mathfrak{m}_{i}(x,y),$ where $\mathfrak{m}_{0}=\mathfrak{m}$ and the $\mathfrak{m}_{i}\in\mathfrak{D}^{1}(L)$ satisfy the condition $[\mathfrak{m}_{t},\mathfrak{m}_{t}]_{MN}=0$.
@@
Moreover, since $[\mathfrak{m}_{t},\mathfrak{m}_{t}]_{MN}=0$, it corresponds to a left symmetric Lie Rinehart algebra structure. In particular, it yields a $t$-parameterized family of products $\mathfrak{m}_{t}:L\otimes L\to L$ and a family of maps $\ell_{t}:L\rightarrow Der(A),$ which satisfy the following identities for all $x,y\in L$:
@@
$\ell_{t}(x)=$ $\ell(x)+\displaystyle\sum_{i\geq 1}t^{i}\Xi_{\mathfrak{m}_{i}}(x).$
The $t$-parametrized family $(L,A,\mathfrak{m}_{t},\ell_{t})$ is called a $1$-parameter formal deformation of $(L,A,\mathfrak{m},\ell)$ generated by $\mathfrak{m}_{1},\cdots,\mathfrak{m}_{m}\in\mathfrak{D}^{1}(L)$.
@@
$\mathfrak{m}_{t}(\mathfrak{m}_{t}(x,y),z)-\mathfrak{m}_{t}(x,\mathfrak{m}_{t}(y,z))=\mathfrak{m}_{t}(\mathfrak{m}_{t}(y,x),z)-\mathfrak{m}_{t}(y,\mathfrak{m}_{t}(x,z)).$ (16)
$\mathfrak{m}_{t}(ax,y)=a\mathfrak{m}_{t}(x,y),$ (17)
$\mathfrak{m}_{t}(x,ay)=a\mathfrak{m}_{t}(x,y)+\ell_{t}(x)ay$ (18)
The identities (17)–(18), mean that $\mathfrak{m}_{i}\in\mathfrak{D}^{1}(L)$. Comparing the coefficients of $t^{n}$ for $n\geq 0$ in equation (16), we get the following:
$\sum_{i+j=n}\mathfrak{m}_{i}(\mathfrak{m}_{j}(x,y),z)-\mathfrak{m}_{i}(x,\mathfrak{m}_{j}(y,z))-\mathfrak{m}_{i}(\mathfrak{m}_{j}(y,x),z)+\mathfrak{m}_{i}(y,\mathfrak{m}_{j}(x,z))=0.$ (19)
For $n=1$, equation (19) implies
@@
The $1$-degree multiderivation $\mathfrak{m}_{1}$ is called the infinitesimal of the deformation $\mathfrak{m}_{t}$. More generally, if $\mathfrak{m}_{i}=0$ for $11\leq i\leq n-$ 1 and $\mathfrak{m}_{n}$ is non zero $1$-degree multiderivation then $\mathfrak{m}_{n}$ is called the $n$-infinitesimal of the deformation $\mathfrak{m}_{t}$. By the above discussion, the following proposition follows immediately.
@@
Now we give a notion of equivalence of two deformations. Let us denote a deformation $(L,A,\mathfrak{m}_{t},\ell_{t})$ of $(L,A,\mathfrak{m},\ell)$ simply by $L_{t}$. Let us consider two deformations $L_{t}$ and $L^{\prime}_{t}$ of $(L,A,\mathfrak{m},\ell)$, generated by $\mathfrak{m}_{i}$ and $\mathfrak{m}^{\prime}_{i}$, respectively, for $i\geq 0$.
@@
$\Phi_{t}:L[[t]]\rightarrow L[[t]]\text{ defined as }\Phi_{t}=id_{L}+\sum_{i\geq 1}t^{i}\phi_{i}$
@@
Proof 6.6.
Let $\Phi_{t}$ be an equivalence of deformation between $\mathfrak{m}_{t}$ and $\tilde{\mathfrak{m}}_{t}$. Then we get,
@@
$\tilde{\mathfrak{m}}_{1}(x,y)+\Phi_{1}(\mathfrak{m}_{0}(x,y))=\mathfrak{m}_{1}(x,y)+\mathfrak{m}_{0}(\Phi_{1}(x),y)+\mathfrak{m}_{0}(x,\Phi_{1}(y)),$
@@
This establishes the result.
@@
Proof 6.9.
Let $\mathfrak{m}_{t}$ be a deformation of left-symmetric Rinehart algebra with n-infinitesimal $\mathfrak{m}_{n}$ for some $n\geq 1$. Assume that there exists a 2-cochain $\phi\in C^{1}(L,L)$ with $\delta(\phi)=\mathfrak{m}_{n}$. Then set
@@
So, ${\bar{\mathfrak{m}}_{n}}=0$. We can repeat the argument to kill off any infinitesimal, which is a coboundary.
@@ 6.2 Obstructions to the extension theory of deformations
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra. Now we consider the problem of extending a deformation of $\mathfrak{m}$ of order $n$ to a deformation of $\mathfrak{m}$ of order $(n+1)$. Let $\mathfrak{m}_{t}$ and $\ell_{t}$ be a deformation of order $n$ of $\mathfrak{m}$ and $\ell$ respectively. That is
@@
where $\mathfrak{m}_{i}\in\mathfrak{D}^{1}(L)$ and $\ell_{i}:L\to Der(A)$ a linear map for each $1\leq i\leq n$ such that
$\mathfrak{m}_{i}(\mathfrak{m}_{j}(x,y),z)-\mathfrak{m}_{i}(x,\mathfrak{m}_{j}(y,z))=\mathfrak{m}_{i}(\mathfrak{m}_{j}(y,x),z)-\mathfrak{m}_{i}(y,\mathfrak{m}_{j}(x,z)),$ (20)
$\mathfrak{m}_{i}(ax,y)=a\mathfrak{m}_{i}(x,y),$ (21)
$\mathfrak{m}_{i}(x,ay)=a\mathfrak{m}_{i}(x,y)+\ell_{i}(x)ay$ (22)
for all $1\leq i,j\leq n$. If there exists a 2-cochain $\mathfrak{m}_{n+1}\in\mathfrak{D}^{1}(L)$ and $\ell_{n+1}:L\to Der(A)$ such that $(L,A,\tilde{\mathfrak{m}_{t}},\tilde{\ell_{t}})$ is a deformation of $(L,A,\mathfrak{m},\ell)$ of order $n+1$, where
@@
Let $\mathfrak{m}_{t}$ be a deformation of $\mathfrak{m}$ of order $n$. Consider the cochain in $C^{3}(L,L)$ defined as
$\begin{split}Obs_{L}(x,y,z)=\sum_{\begin{subarray}{c}i+j=n+1;\\ i,j>0\end{subarray}}\Big{(}\mathfrak{m}_{i}(\mathfrak{m}_{j}(x,y),z)-\mathfrak{m}_{i}&(x,\mathfrak{m}_{j}(y,z))\\ &-\mathfrak{m}_{i}(\mathfrak{m}_{j}(y,x),z)+\mathfrak{m}_{i}(y,\mathfrak{m}_{j}(x,z)))\Big{)},\end{split}$ (23)
@@
Proof 6.14.
Suppose that a deformation $\mathfrak{m}_{t}$ of order $n$ extends to a deformation of order $n+1$. Then
$\sum_{\begin{subarray}{c}i+j=n+1;\\ i,j\geq 0\end{subarray}}\Big{(}\mathfrak{m}_{i}(\mathfrak{m}_{j}(x,y),z)-\mathfrak{m}_{i}(x,\mathfrak{m}_{j}(y,z))-\mathfrak{m}_{i}(\mathfrak{m}_{j}(y,x),z)+\mathfrak{m}_{i}(y,\mathfrak{m}_{j}(x,z)))\Big{)}=0.$
@@
$Obs_{L}=\delta(\mathfrak{m}_{n+1})$
for some 2-cochain $\mathfrak{m}_{n+1}$. Define a map $\tilde{\mathfrak{m}_{t}}:L[[t]]\times L[[t]]\rightarrow L[[t]]$ as follows
$\tilde{\mathfrak{m}_{t}}=\mathfrak{m}_{t}+\mathfrak{m}_{n+1}t^{n+1}.$
@@
$\sum_{\begin{subarray}{c}i+j=n+1;\\ i,j\geq 0\end{subarray}}\Big{(}\mathfrak{m}_{i}(\mathfrak{m}_{j}(x,y),z)-\mathfrak{m}_{i}(x,\mathfrak{m}_{j}(y,z))-\mathfrak{m}_{i}(\mathfrak{m}_{j}(y,x),z)+\mathfrak{m}_{i}(y,\mathfrak{m}_{j}(x,z)))\Big{)}=0.$
@@
If $H^{3}(L,L)=0,$ then every 2-cocycle in $C^{2}(L,L)$ is the infinitesimal of some deformation of $\mathfrak{m}$.